B15-13 (book/static) : Question Help O (EBIT-EPS analysis) Three recent graduates of the computer science program at the UnivPlease help me with the questions for all the above. Thank you.
Answer #1

a. No of shares with plan A = 2,000,000 / 20 = 100,000 shares

No of shares as per Plan B = 1,000,000 / 20 = 50,000 shares

Interest on the bond issue = 1,000,000 *11% = 110,000

Equating both EBIT's:

( EBIT ) * ( 1 -30%) / 100,000 = ( EBIT - 110,000) (1-0.3) / 50,000

EBIT / 2 = EBIT - 110,000

EBIT = 220,000

The EBIT indifference level associated with the two financing plans is $220,000

b. EPS under Plan A = ( 300,000) * (1-30%) / 100,000 = 2.1

EPS under Plan B = ( 300,000 -110,000) * (1-30%) / 50,000 = 2.66

Plan B will generate a higher EPS.